OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through point (1, -2) and has a slope of 1/3. A. 3x - y - 7 = 0 B. x - 3y + 7 = 0 C. x - 3y - 7 = 0

OpenStudy (anonymous):

The firs thing to do is to set all of the equations into y=mx+b form

OpenStudy (anonymous):

u mean gradient as slope right =), so troll

OpenStudy (anonymous):

y=3x+7 y=1/3x +7/3 y=1/3x+7/3

OpenStudy (anonymous):

We can see that only two of the answers have a 1/3 slope, so the first choice can be eliminated.

OpenStudy (anonymous):

kirby, u sure plus?

OpenStudy (anonymous):

Wait hold on, the last choice is y=1/3x-7/3

OpenStudy (anonymous):

Then plug 1 and -2 in to see which one equals itself

OpenStudy (anonymous):

is C , why 1/3+7/3=-2 one? lol so troll, i must really take my hat off for u if i have a hat, wait here, imma gonna buy one, just saying

OpenStudy (mathmale):

Why not use the point-slope formula for a straight line? It's by far the most straightforward approach. Here it is: y-b = m(x-1), where (a,b) is the point in question and m is the slope.